OpenStudy (anonymous):

f(x)=3cosx−cos^(3)x for 0

OpenStudy (shayaan_mustafa):

Hi jinnie :) for x-intercept put f(x)=0 and solve for x.. That will be your answer.. :) Good luck dear.. :)

OpenStudy (anonymous):

i tried but im not getting the correct ans

OpenStudy (shayaan_mustafa):

don't worry. never give up.. :) math didn't create the human being. but human being created math. so just believe yourself and follow rules.. can you try that solution here ?

OpenStudy (anonymous):

try factoring first

OpenStudy (anonymous):

\[f(x)=3\cos x−\cos^3 x = \cos x (3 - \cos^2(x)) \] One of the actors above can be zero. Which one? Set it to zero and solve

OpenStudy (anonymous):

factors not actors above

OpenStudy (shayaan_mustafa):

lol :D

OpenStudy (anonymous):

yeah it still isnt working when it set it equal to zero